# Long-Context Generalization with Sparse Attention

### v2

@@ Appendix D Non-Dispersion of $\alpha$-entmax
[Threshold Behavior for Two-level Logits] Consider logits $\bm{z}\in\mathbb{R}^{n}$ where $k$ tokens have value $M$ and $(n-k)$ tokens have value $m$ with $M>m$.
@@
[…]
@@
$\operatorname{softmax}(\bm{z}/\theta)_{i}=\frac{\exp(M/\theta)}{k\exp(M/\theta)+(n-k)\exp(m/\theta)}.$ (74)
@@
Part (iii) - Strong concentration resilience of $\alpha$-entmax: First, we focus on the two-level case from Lemma D, where $k$ tokens have logit value $M$ and $(n-k)$ tokens have value $m$. When $\Delta=M-m\geq\frac{k^{-(\alpha-1)}}{\alpha-1}$, only the $k$ tokens with value $M$ receive non-zero attention:
@@ E.1 No Positional Encoding (NoPE)
The first assumption tells that key and query weight matrices are bounded. The second assumption ensures that the node representations’ trajectories across $t$ layers stays within a fixed interval $[-C^{2},C^{2}]$.
@@
For softmax, Wu et al. (2025) have shown that lim_ℓ→∞ ~p^(ℓ)_soft,i1 = 1 for all $1<i\leq n$.
@@
    Edge deletion: For every layer $\ell$, an edge $(j,i)\in\mathcal{G}$ is present in the dynamic graph $\mathcal{G}^{(\ell)}_{\alpha}$ if and only if $(\alpha-1)z^{(\ell)}_{ij}>\tau(\bm{z}^{(\ell)}_{i})$, where $z^{(\ell)}_{ij}=\langle\bm{q}^{(\ell)}_{i},\bm{k}^{(\ell)}_{j}\rangle$ and $\tau(\bm{z}^{(\ell)}_{i})$ is the entmax threshold. Otherwise, $p^{(\ell)}_{\alpha,ij}=0$ and the edge is removed for that layer.
@@
This means the relative behavior of attention weights in $\alpha$-entmax compared to softmax depends on the specific configuration of logits and thresholds, rather than following a simple universal relationship. Since tokens that remain connected must distribute probability mass among fewer options (due to pruning), there exists some $10<\delta_{ij}<$ 1 such that:
@@
$10\leq\lim_{\ell\to\infty}\tilde{p}^{(\ell)}_{\alpha,i1}<$ 1 . (104)
@@ E.2 ALiBi
For a token $j$ with distance $(i-j)$, even with the maximum logit $z^{(h)}_{\max}$:
@@
This proposition establishes that with $\alpha$-entmax and ALiBi positional bias, there exists a head-dependent hard cutoff distance $d^{(h)}_{\max}$ beyond which tokens receive exactly zero attention. This creates an adaptive but bounded attention window that depends on both content relevance ($z^{(h)}_{\max}-z^{(h)}_{\min}$) and the sparsity parameter $\alpha$, naturally limiting the effective context without requiring explicit truncation. This property allows the model to focus computational resources on a relevant window of tokens, which can be particularly valuable for efficiently processing long documents.
@@ E.3 RoPE
Here, we simplified the last step by focusing on the quadratic decay from the cosine term while omitting the sine terms $-\sin(\phi_{ijk})g_{k}(j-i)$. For semantically aligned tokens where $\phi_{ijk}\approx 0$, the sine term’s contribution is minimal since $\lim_{x\to 0}\sin(x)=0$.
@@ E.4 Comparison between Positional Encoding Methods with $\alpha$-entmax
These interactions between positional encodings and $\alpha$-entmax have important practical implications, and further motivate the introduction of our hybrid approach, NAPE (§4.2). By creating natural, content-adaptive attention windows, NAPE combine the benefits of sparse, focused attention with awareness of token positions, allowing models to effectively balance local and global information processing. This provides a principled alternative to manually designed sparse attention patterns like sliding windows or dilated attention, with the advantage of adapting to content relevance rather than using fixed patterns.
@@ G.1 Synthetic Data
Following the data diversity assumptions of Zhou et al. (2024), we generate a large number of samples—between 10 million and 50 million, depending on task complexity. See Table 2 for training/test details on for each task along with model hyperparameters. The 2Back and Local Count tasks are token-classification tasks, while the remaining tasks are generative. In Figure 10, we show examples of the tasks we introduce in this work.
@@
Flip-Flop uses sequences made up of alternating instructions "write", "ignore", and "read" (w, i ,r). Each instruction is paired with a bit (0 or 1). The model must memorize the bit if it follows a w, and recall the last stored bit when it encounters an r. For example, in the string "i 0 w 0 i 1 i 1 w 1 i 0 i 1 r", the correct output is 1. We conducted experiments on two datasets with different “write” instruction probabilities: $10\%$ sparse, $80\%$ dense.
@@
MQMTAR is a retrieval task in which models must produce a sequence of multi-token values corresponding to the queries provided at the end of the input sequence (see Figure 10). MQMTAR employs three special tokens: (1) 0 for empty space; (2) 1 as the key–value delimiter; and (3) 3 as the query delimiter, which is also used to separate values in the target sequence. We set the lengths of both keys and values to be 2 tokens, resulting in 5 tokens per key-value pair in the input. The number of queries is 4, and the density of key-value pairs is $80\%$ of the total number of tokens. Finally, the size of alphabet is 256 from which we constructed 100K key-value pairs.
@@
G.2 Models
All synthetic tasks are trained with a decoder-like transformer. We evaluate models in extreme settings by using as few layers as possible as our aim is to test the attention mechanism coupled with our positional-encoding strategy, rather than the scaling capabilities of transformer. However, for the Reverse task—which proved particularly challenging for softmax-based models—we increment the layer count until the softmax baseline generalizes to at least 1.5x the in-distribution length.
For experiments with RoPE, we use the Hugging Face implementation from LLaMA 3 (Grattafiori et al., 2024), which includes RoPE scaling. Since our sequences are relatively short, the base frequency is set to its default value of $10{,}000$. To improve length extrapolation in RoPE-based models, we apply a scaling factor of $16$, which we found to be optimal for Flip-Flop under 4×4$\times$4 × extrapolation (see Table H.6); factors of $8$ or $32$ degrade performance. For NAPE, each ALiBi head uses a slope of $m=\frac{1}{h}$, where $h$ is the head index. We employ 8 attention heads for all tasks except Copy and MQMTAR, where 16 heads yields a performance boost across all models.
In case of ASEntmax, $\beta^{i}_{h}$ and $\gamma^{i}_{h}$ are computed per token $i$ via linear projections followed by activations softplus and tanh respectively, allowing to scale attention adaptively based on content. For our experiments with $\alpha$-entmax, we use $\alpha=1.5$ as the default value for the $\alpha$-entmax models, unless mentioned otherwise. Furthermore, we use the Gemma2 implementation from Hugging Face, but disabling sliding-window attention in all layers. For experiments with $\alpha$-entmax, we replaced FlashAttention with AdaSplash (Gonçalves et al., 2025).
@@ G.2 Models
We observe that even when models are $100\%$ accurate in-distribution, they still require significantly more training; in some cases, the training loss reached as low as $10^{-8}$. Therefore, the best checkpoint is selected based on performance at 8×8$\times$8 × the in-distribution sequence length. In some cases such as the Sort task and models with RoPE, where generalization up to 8×8$\times$8 × was not possible, we use BLEU as an intermediate metric and 2×2$\times$2 × or 4×4$\times$4 × the in-distribution sequence length. We perform evaluation with $1K$ samples per sequence length. 2Back and Local Count are evaluated using accuracy, as they are classification tasks, whereas for the remaining tasks, we use exact match accuracy—assigning 1 only if the entire predicted sequence matches the reference and 0 otherwise. We report results for the single best-performing model selected from experiments conducted with multiple random seeds (3) and various learning rates.
Finally, for some tasks, we also report results for SEntmax—which learns $\beta$ and $\gamma$ directly, without linear projections or nonlinearities—and also for ASSMax—which applies our adaptive-scaling strategy to Softmax.
@@
Appendix H Additional Results

### v3

@@ Appendix C Non-Dispersion of $\alpha$-entmax
[Threshold Behavior for Two-level Logits] Consider logits $\bm{z}\in\mathbb{R}^{n}$ where $k$ tokens have value $M$ and (nk)(n-k) tokens have value $m$ with $M>m$.
@@
Thus, when $\Delta\geq\frac{k^{-(\alpha-1)}}{\alpha-1}$, only the $k$ tokens with value $M$ receive non-zero attention, each receiving an attention of $\frac{1}{k}$. Therefore, when $\Delta\geq\frac{k^{-(\alpha-1)}}{\alpha-1}$, the attention weights for $\alpha\text{-entmax}$ are:
@@
$\mathrm{softmax}(\bm{z}/\theta)_{i}=\frac{\exp(M/\theta)}{k\exp(M/\theta)+(n-k)\exp(m/\theta)}.$ (29)
@@
Part (iii) - Strong concentration resilience of $\alpha$-entmax: First, we focus on the two-level case from Lemma C, where $k$ tokens have logit value $M$ and (nk)(n-k) tokens have value $m$. When $\Delta=M-m\geq\frac{k^{-(\alpha-1)}}{\alpha-1}$, only the $k$ tokens with value $M$ receive non-zero attention:
@@
Appendix D Representational Collapse and Over-squashing
D.1 Proof of Lemma 3.1
For $\alpha>1$, we need to find the value $b_{\max}$ such that, for any $b\leq b_{\max}$, $\operatorname{\alpha\text{-}entmax}(\bm{x}^{*})_{n}=0$ holds. From the definition of $\alpha$-entmax equation 2, a token $i$ receives non-zero probability iff $(\alpha-1)z_{i}>\tau(\bm{z})$, where $\tau(\bm{z})$ is the threshold ensuring the sum of probabilities equals 1. Therefore, for the token $b$ in the extended sequence $\bm{x}^{*}$ to receive zero probability (thus not affecting other probabilities), we need:
$(\alpha-1)b\leq\tau(\bm{x}^{*}).$ (47)
$(\alpha-1)b\leq\tau(\bm{x}).$ (48)
Solving for $b$, we get:
$b\leq\frac{\tau(\bm{x})}{\alpha-1}.$ (49)
Thus, we can define $b_{\max}=\frac{\tau(\bm{x})}{\alpha-1}$. For any $b\leq b_{\max}$, the token at position $n$ in $\bm{x}^{*}$ receives zero attention, meaning it doesn’t affect the normalization. Therefore, $\tau(\bm{x}^{*})=\tau(\bm{x})$, which means that the condition equation 49 is both necessary and sufficient, and:
D.2 Proof of Proposition 3.3 for Representational Preservation
We prove the first part of Proposition 3.3 by exhibiting the counterexample below, following the synthetic construction introduced by Barbero et al. (2024).
{proposition}
where the final token $\bm{v}_{a}\in\mathbb{R}^{d}$ is repeated.
$\|\bm{v}_{n}^{(L)}-\bm{v}_{n+1}^{*(L)}\|_{1}\geq c>0$ (52)
Proof.
We prove this through explicit construction.
Consider the specific embedding choice where:
This construction yields the following attention logits:
$\bm{z}_{n}^{(1)}$ $=[b,b,\ldots,b,\phi],$ (55)
$\bm{z}_{n+1}^{*(1)}$ $=[b,b,\ldots,b,\phi,\phi].$ (56)
Specific counterexample.
$\bm{z}_{n}^{(1)}$ $=[0,\ldots,0,0.5],$ (57)
$\bm{z}_{n+1}^{*(1)}$ $=[0,0,\ldots,0,0.5,0.5].$ (58)
For $\alpha=2.0$ (sparsemax), the attention distributions are:
$\text{sparsemax}(\bm{z}_{n}^{(1)})$ $=[p_{b},p_{b},\ldots,p_{b},p_{n}]\quad\text{(dense)}$ (59)
$\text{sparsemax}(\bm{z}_{n+1}^{*(1)})$ $=[0,0,\ldots,0,\frac{1}{2},\frac{1}{2}]\quad\text{(sparse)}$ (60)
where $0<p_{b}<p_{n}<1$. As $n\to\infty$, we have $p_{n}\to p^{*}$ where:
This demonstrates that the $L_{1}$ difference remains bounded at approximately $c=0.354$, independent of sequence length $n$. This specific example already establishes the existence of a counterexample to representational collapse with $\alpha$-entmax. We now extend this result to prove Proposition D.2 for general constructions of $\bm{v}$ and $\bm{v}_{a}$.
General Construction.
With values of $\phi$ such that $b+2^{-(\alpha-1)}/(\alpha-1)\leq\phi<b+1/(\alpha-1)$, $\alpha$-entmax produces the following distributions: [^4] The upper bound ensures a dense output for $\bm{z}_{n}^{(1)}$, following Lemma 3.1 with $\tau=(\alpha-1)\phi-1$. The lower bounds ensure a sparse output for $\bm{z}_{n+1}^{*(1)}$, following Lemma C with $k=2$.
In particular, since the first n1n-1 positions share the same representation, we have $p^{(1)}_{1}=p^{(1)}_{2}=...=p^{(1)}_{n-1}=(1-p^{(1)}_{n})/(n-1)=p^{(1)}_{b}>0$, with $0<p^{(1)}_{n}<1$. This leads to representations:
$\bm{v}_{n+1}^{*(1)}$ $=\frac{1}{2}\bm{v}_{a}+\frac{1}{2}\bm{v}_{a}=\bm{v}_{a}.$ (66)
$=(1-p_{n}^{(1)})\left\|\bar{\bm{v}}-\bm{v}_{a}\right\|_{1}.$ (68)
First condition.
We need to choose parameters so that as $n\to\infty$, the original sequences remains dense and the extended sequence is in the sparse regime. From our analysis with Lemma 3.1 and C, this requires:
$\frac{2^{-(\alpha-1)}}{\alpha-1}\leq\phi-b<\frac{1}{\alpha-1}.$ (69)
From the $\alpha$-entmax definition, we have:
$(p_{b}^{(1)})^{\alpha-1}$ $=(\alpha-1)b-\tau,$ (70)
$(p_{n}^{(1)})^{\alpha-1}$ $=(\alpha-1)\phi-\tau,$ (71)
From the normalization constraint $(n-1)p_{b}^{(1)}+p_{n}^{(1)}=1$:
$p_{b}^{(1)}=\frac{1-p_{n}^{(1)}}{n-1}.$ (73)
Substituting:
We know from the dense regime that $p_{n}^{(1)}\to p^{*}$, with $0<p^{*}<1$. Thus, as $n\to\infty$, the probability $p_{n}^{(1)}$ satisfies:
$(p^{*})^{\alpha-1}=(\alpha-1)(\phi-b)$ (76)
Therefore:
$p^{*}=[(\alpha-1)(\phi-b)]^{\frac{1}{\alpha-1}}<1.$ (77)
Second condition.
  • $\bm{v}_{i}=\frac{b}{\sqrt{\phi}}\bm{e}_{1}+\bm{e}_{2}$ for $i=1,\ldots,n-1$.
Note that this construction satisfies the logit constraints:
$\bm{v}_{a}^{T}\bm{v}_{i}$ $=\sqrt{\phi}\cdot\frac{b}{\sqrt{\phi}}+0\cdot 1=b,$ (78)
$\bm{v}_{a}^{T}\bm{v}_{a}$ $=(\sqrt{\phi})^{2}=\phi.$ (79)
The average representation is:
Since $\bm{v}_{a}=\sqrt{\phi}\bm{e}_{1}$, we have:
The bound is strictly positive because $b\neq\phi$ by the logit difference requirement, and because of the constant +1+1 term from the $\bm{e}_{2}$ component. Therefore, $\lim_{n\to\infty}\|\bar{\bm{v}}-\bm{v}_{a}\|_{1}\not\to 0$. For the case $d=1$, we can use the simpler construction $v_{i}=\frac{b}{\sqrt{\phi}}$ and $v_{a}=\sqrt{\phi}$, where the non-collapse condition becomes verifying that $\frac{b}{\sqrt{\phi}}\neq\sqrt{\phi}$, which follows from $b\neq\phi$. In contrast, as shown by Barbero et al. (2024), the resulting representations become increasingly similar as $n\to\infty$ with softmax ($\alpha=1.0$), regardless of the input content, leading to representational collapse. ∎
Empirical Verification of Representational Preservation.
  1. 1.
    Constant prefixes: $b=1$ and $\phi=1.2$.
  2. 2.
    Random prefixes: $b\sim\mathcal{U}(0,1)$ and $\phi=1.2$.
D.3 Proof of Proposition 3.3 for Over-squashing Alleviation
{proposition}
$\bm{y}_{n}=\mathrm{norm}_{3}\bigl(\bm{v}_{n}^{(L)}\bigr)$ (82)
Proof.
[…]
$\mathcal{T}_{\ell}(k_{\ell-1})$ $=\{k\in\{k_{\ell-1},k_{\ell-1}+1,\ldots,n\}:k\in\mathcal{S}_{k_{\ell-1}}^{(\ell-1)}\text{ or }k=k_{\ell-1}\}\quad\text{for }\ell>1.$ (90)
These sets have the following meaning:
Empirical Verification of Over-squashing Alleviation.
@@ E.1 No Positional Encoding (NoPE)
The first assumption tells that key and query weight matrices are bounded. The second assumption ensures that the node representations’ trajectories across $t$ layers stays within a fixed interval $C^{2},C^{2}]$.
@@
For softmax, Wu et al. (2025) have shown that
$\lim_{\ell\to\infty}\tilde{p}^{(\ell)}_{\text{soft},i1}=1$
for all $1<i\leq n$.
@@
[…]
@@
This means the relative behavior of attention weights in $\alpha$-entmax compared to softmax depends on the specific configuration of logits and thresholds, rather than following a simple universal relationship. Since tokens that remain connected must distribute probability mass among fewer options (due to pruning), there exists some $0<\delta_{ij}<1$ such that:
@@
$0\leq\lim_{\ell\to\infty}\tilde{p}^{(\ell)}_{\alpha,i1}<1.$ (103)
@@ E.2 ALiBi
For a token $j$ with distance (ij)(i-j), even with the maximum logit $z^{(h)}_{\max}$:
@@
[…]
@@ E.3 RoPE
Here, we simplified the last step by focusing on the quadratic decay from the cosine term while omitting the sine terms $\sin(\phi_{ijk})g_{k}(j-i)$. For semantically aligned tokens where $\phi_{ijk}\approx 0$, the sine term’s contribution is minimal since $\lim_{x\to 0}\sin(x)=0$.
@@ E.4 Comparison between Positional Encoding Methods with $\alpha$-entmax
These interactions between positional encodings and $\alpha$-entmax have important practical implications, and further motivate the introduction of our hybrid approach, NAPE (NoPE + ALiBi). By creating natural, content-adaptive attention windows, NAPE combine the benefits of sparse, focused attention with awareness of token positions, allowing models to effectively balance local and global information processing. This provides a principled alternative to manually designed sparse attention patterns like sliding windows or dilated attention, with the advantage of adapting to content relevance rather than using fixed patterns.
@@ G.1 Synthetic Data
Following the data diversity assumptions of Zhou et al. (2024), we generate a large number of samples—between 10 million and 50 million, depending on task complexity. See Table 5 for training/test details on for each task along with model hyperparameters. The 2Back and Local Count tasks are token‐classification tasks, while the remaining tasks are generative. In Figure 8, we show examples of the tasks we introduce in this work.
@@
Flip-Flop uses sequences made up of alternating instructions "write", "ignore", and "read" (w, i ,r). Each instruction is paired with a bit (0 or 1). The model must memorize the bit if it follows a w, and recall the last stored bit when it encounters an r. For example, in the string "i 0 w 0 i 1 i 1 w 1 i 0 i 1 r", the correct output is 1. We conducted experiments on two datasets with different “write” instruction probabilities: $10\%$ - sparse, $80\%$ - dense.
@@
MQMTAR is a retrieval task in which models must produce a sequence of multi‐token values corresponding to the queries provided at the end of the input sequence (see Figure 8). MQMTAR employs three special tokens: (1) 0 for empty space; (2) 1 as the key–value delimiter; and (3) 3 as the query delimiter, which is also used to separate values in the target sequence. We set the lengths of both keys and values to be 2 tokens, resulting in 5 tokens per key-value pair in the input. The number of queries is 4, and the density of key-value pairs is $80\%$ of the total number of tokens. Finally, the size of alphabet is 256 from which we constructed 100K key-value pairs.
@@
G.2 Models for Synthetic Tasks
All synthetic tasks are trained with a decoder-like transformer. We evaluate models in extreme settings by using as few layers as possible as our aim is to test the attention mechanism coupled with our positional‐encoding strategy, rather than the scaling capabilities of transformer. However, for the Reverse task—which proved particularly challenging for softmax‐based models—we increment the layer count until the softmax baseline generalizes to at least 1.5x the in-distribution length.
For experiments with RoPE, we use the Hugging Face implementation from LLaMA 3 (Grattafiori et al., 2024), which includes RoPE scaling. Since our sequences are relatively short, the base frequency is set to its default value of 10,00010{,}000. To improve length extrapolation in RoPE‐based models, we apply a scaling factor of $16$, which we found to be optimal for Flip‐Flop under $4\times$ extrapolation (see Table H.6); factors of $8$ or $32$ degrade performance. For NAPE, each ALiBi head uses a slope of $m=\frac{1}{h}$, where $h$ is the head index. We employ 8 attention heads for all tasks except Copy and MQMTAR, where 16 heads yields a performance boost across all models.
In case of ASEntmax, $\beta^{i}_{h}$ and $\gamma^{i}_{h}$ are computed per token $i$ via linear projections followed by activations softplus and tanh respectively, allowing to scale attention adaptively based on content. For our experiments with $\alpha$-entmax, we use $\alpha=1.5$ as the default value for the $\alpha$-entmax models, unless mentioned otherwise. Furthermore, we use the Gemma2 implementation from Hugging Face, but disabling sliding‐window attention in all layers. For experiments with $\alpha$-entmax, we replaced FlashAttention with AdaSplash (Gonçalves et al., 2025).
@@ G.2 Models for Synthetic Tasks
We observe that even when models are $100\%$ accurate in-distribution, they still require significantly more training; in some cases, the training loss reached as low as $10^{-8}$. Therefore, the best checkpoint is selected based on performance at $8\times$ the in-distribution sequence length. In some cases such as the Sort task and models with RoPE, where generalization up to $8\times$ was not possible, we use BLEU as an intermediate metric and $2\times$ or $4\times$ the in-distribution sequence length. We perform evaluation with $1K$ samples per sequence length. 2Back and Local Count are evaluated using accuracy, as they are classification tasks, whereas for the remaining tasks, we use exact match accuracy—assigning 1 only if the entire predicted sequence matches the reference and 0 otherwise. We report results for the single best-performing model selected from experiments conducted with multiple random seeds (3) and various learning rates.
For some tasks, we also report results for SEntmax—which learns $\beta$ and $\gamma$ directly, without linear projections or nonlinearities—and also for ASSMax—which applies our adaptive-scaling strategy to Softmax. Finally, for some tasks we also experiment with Stick-Breaking Attention (SB, Tan et al. 2025), which corresponds to a different positional encoding strategy.
@@
G.3 Language Modeling
We train 420M-parameter decoder-only models following the LLaMA-3 architecture, using the following hyperparameters: (1) model dimension: 768; (2) number of layers: 24; (3) number of heads: 12. For RoPE and FFN layers, we use the default LLaMA-3 settings. The training is conducted using the torchtitan library (Liang et al., 2025). We train for a total of 13.5K steps, with 1.35K steps allocated for warm-up. The learning rate is set to $3\cdot 10^{-4}$, and the cooldown phase corresponds to $10\%$ of the total steps with a minimum learning rate of $3\cdot 10^{-5}$, where we apply cosine decay scheduling. We use the AdamW optimizer and train with bfloat16 precision. The global batch size is set to 524K tokens. For tokenization, we employ the LLaMA-3.2-1B tokenizer. For NAPE experiments, we choose original exponential ALiBi slopes.
We use the following benchmarks for short context evaluation: Lambada (Paperno et al., 2016), HellaSwag (Zellers et al., 2019), PIQA (Bisk et al., 2020), ARC-C (Clark et al., 2018), Winogrande (Sakaguchi et al., 2020), and OpenBookQA (Mihaylov et al., 2018). For length generalization, we use 500 documents from each dataset (ArXiv and PubMed), with all documents at least 16K tokens long, and 500 samples (default number of samples in RULER) for each length for S-NIAH-1 and S-NIAH-2. We evaluate models with RoPE in two settings: vanilla RoPE and RoPE with Adjusted Base Frequency (ABF) (Xiong et al., 2024), where the base frequency is set to 500,000500{,}000, i.e., $50\times$ used in pre-training. Results for language modeling are in Appendix H.9.
Appendix H Detailed Results